# Supervariable Approach

### 16 Feb 2024 v2

@@ 6 Conclusions
In our present endeavour, we have studied the FLRW model of homogeneous and isotropic Universe within the framework of supervariable approach to BRST formalism. We have derived the off-shell nilpotent and absolutely anticommuting (anti-)BRST transformations corresponding to the local time-reparametrization transformation by exploiting the BT supervariable approach where, the property of (super-)diffeomorphism invariance is taken into account. Interestingly, the CF condition, necessary for the proof of absolute anticommutativity of the (anti-)BRST transformations, emerges naturally in this approach. Moreover, using the basic principles of BRST formalism, we have explicitly obtained the gauge-fixed (anti-)BRST invariant Lagrangians ${\cal L}_{(b)}$ and ${\cal L}_{(ab)}$. Both Lagrangians are coupled (but equivalent) because of the existence of CF-type condition (33). Further, a close observation reveals that both Lagrangians as well as (anti-)BRST transformations are symmetric if we interchange the role of $(\bar{C})C$ and $(\bar{B})B$. To be more precise, if we replace $C\to+\bar{C}$, $\bar{C}\to-C$, $B\to-\bar{B}$ one can go from $L_{(b)}\to L_{(ab)}$ and from $s_{b}\to s_{ab}$.

### 05 Mar 2024 v3

@@ 6 Conclusions
In our present endeavour, we have studied the FLRW model of homogeneous and isotropic Universe within the framework of supervariable approach to BRST formalism. We have derived the off-shell nilpotent and absolutely anticommuting (anti-)BRST transformations corresponding to the local time-reparametrization transformation by exploiting the BT supervariable approach where, the property of (super-)diffeomorphism invariance is taken into account. Interestingly, the CF condition, necessary for the proof of absolute anticommutativity of the (anti-)BRST transformations, emerges naturally in this approach. Moreover, using the basic principles of BRST formalism, we have explicitly obtained the gauge-fixed (anti-)BRST invariant Lagrangians $L_{(b)}$ and $L_{(ab)}$. Both Lagrangians are coupled (but equivalent) because of the existence of CF-type condition (33). Further, a close observation reveals that both Lagrangians as well as (anti-)BRST transformations are symmetric if we interchange the role of $(\bar{C})C$ and $(\bar{B})B$. To be more precise, if we replace $C\to+\bar{C}$, $\bar{C}\to-C$, $B\to-\bar{B}$ one can go from $L_{(b)}\to L_{(ab)}$ and from $s_{b}\to s_{ab}$.